find the area of trapezium when parallel sides are 20 m & 30 m and non parallel sides are 6 m & 8m

Grade:8

1 Answers

Saurabh Koranglekar
5 years ago
Let h is the height of trapezium
so,
6^2=h^2+x^2
8^2=h^2+(10-x)^2


64-36=(10-x)^2- x^2
28=(10-x-x) 10
2.8=10-2x
7.2 = 2x
x=3.6 m

and 10-3.6 =6.4 m

now,
h={6^2-(3.6)^2}^1/2=4.8 m

now,
area of trapezium=1/2 (20+30) x 4.8
=2.4 x 50=120 m^2